(6 pts) A particular state computed income tax, T (in dollars) on a person’s income, x (in dollars), by the following formula:
T (x)= {0.018x if 0 ≤ x ≤ 5,000
90 + 0.025 (x-5,000) if 5,000 < x ≤ 10,000
215 + 0.050 (x-10,000) if x > 10,000
What is the tax on an income of $7,300? Show work.



(b) What is the tax on an income of $73,000? Show work.


Answer by CubeyThePenguin(3113)

a) 7300 is between 5000 and 10000.
T(7300) = 90 + 0.025(7300 - 5000) = 147.5

b) 73000 is greater than 10,000.
T(73000) = 215 + 0.05(73000 - 10000) = 3365